OpenStudy (anonymous):

A rectangle garden has a perimeter of 172 feet. The length of the garden Is 24 feet more than the width, what are the dimensions of the garden?

OpenStudy (anonymous):

 p= 172 f=2x+2y...........(1) x=y+24................(2) so substitute eq (2) in (1): 172=2(y+24)+2y 172=4y+48 the width : y=(172-48)/4=31 feet the length:x=31+24=55 feet